# Piecewise Polynomial Tensor Network Quantum Feature Encoding

### v1

@@ Abstract
This work introduces a novel method for embedding continuous variables into quantum circuits via piecewise polynomial features, utilizing low-rank tensor networks. Our approach, termed Piecewise Polynomial Tensor Network Quantum Feature Encoding (PPTNQFE), aims to broaden the applicability of quantum algorithms by incorporating spatially localized representations suited for numerical applications like solving partial differential equations and function regression. We demonstrate the potential of PPTNQFE through efficient point evaluations of solutions of discretized differential equations and in modeling functions with localized features such as jump discontinuities. Although promising, challenges such as unexplored noise impact and design of trainable circuits remain. This study opens new avenues for enhancing quantum models with novel feature embeddings and leveraging TN representations for a wider array of function types in quantum machine learning.
@@ 1.1 Quantum Advantage in Quantum Machine Learning
The rationale for employing PQCs often hinges on the large size of the quantum state space. However, these circuits often produce models similar to Fourier series, which may not fully take advantage of the exponentially large Hilbert space and can be emulated classically Schuld et al. (2021); Kordzanganeh et al. (2023); Shin et al. (2023), or they may be overly expressive, hindering effective training and generalization Kübler* et al. (2021); Huang et al. (2021b); Jerbi et al. (2023); Cerezo et al. (2023). Furthermore, the impact of statistical and hardware-induced errors is frequently overlooked, potentially undermining any claims of enhanced expressiveness or precision.
@@ 1.2 Inductive Bias and Feature Encoding
The efficacy of a model extends beyond mere expressiveness; it depends on the synergy between sufficient expressiveness and the appropriate, problem-specific inductive bias. Recent innovative studies have explored the potential to imbue models with biases that are distinctively quantum in nature Bowles et al. (2023); Gili et al. (2023); Nguyen et al. (2022); Schatzki et al. (2024).
@@ 1.3 This Work
Our approach leverages an intermediate tensor network (TN) representation, which makes it adaptable to any function that can be represented in low-rank format Oseledets (2013); Ali and Nouy (2023). We denote these embeddings as Piecewise Polynomial Tensor Network Quantum Feature Encodings (PPTNQFE). It is crucial to note that these embeddings are not inherently superior or inferior to others and do not confer any QA. Through toy examples, we elucidate the scenarios where PPTNQFEs are advantageous and when they are not.
The structure of this paper is as follows. Section 2 reviews piecewise polynomials and the properties required for embedding. Section 3 discusses TNs and functional TN representations. Section 4 details the methodology for expressing piecewise polynomial bases as low-rank TNs and their subsequent translation into quantum circuits. Section 5 presents two important applications: point evaluations of PDE solutions and function learning (regression). The paper concludes in Section 6. The implementation and examples provided are confined to one spatial dimension, with further details in Section 7. While extending these methods efficiently to higher dimensions is theoretically feasible, it would entail a more involved implementation.
@@ 2 Piecewise Polynomials
A plot of the hat functions is presented in Figure 1. These satisfy conditions (I), (II) and (III).
Figure 1: 1D piecewise linear hat basis for a grid of 8 uniformly spaced points, including the boundary $x\in\{\pm 1\}$
@@ 3 Tensor Networks
Tensor networks are an implicit, parsimonious representation of high-order tensors. They are most commonly known for their effectiveness in modeling high-dimensional quantum systems with low entanglement. For details on TN formats and diagrams, see Biamonte and Bergholm (2017). In this work, we will be using the matrix product state (MPS) format Schollwöck (2011), also known under the name tensor train (TT) format Oseledets (2011), see Figure 2.
@@ 4 Methodology
The case of even $k$ is similar. Since the maximum rank is $2$, converting to a quantum circuit as in Figure 2. we require only $2$-qubit gates.
@@ 5 Applications
In this section, we dive into two illustrative applications of the PPTNQFE method that we have introduced. The first application focuses on the efficient point evaluation of solutions to differential equations. In this scenario, the discretized and normalized solution is encoded in a quantum state, with the basis coefficients represented within the quantum amplitudes. The second application we explore is regression, specifically the supervised learning of functions depending on continuous variables from given data samples. Here, PPTNQFE allows, e.g., the approximation of jump functions in a particularly simple manner.
@@ 5.1 Differential Equations
The associated linear system, represented in matrix form,
$\boldsymbol{A}\boldsymbol{u}=\frac{1}{h}\begin{pmatrix}2&-1&0&\dots&\\ -1&2&-1&\dots&\\ &&\ddots&&\\ &&&&-1\\ &&&-1&2\end{pmatrix}\boldsymbol{u}=\boldsymbol{f}.$
@@ 5.2 Regression
In the context of supervised learning, we are presented with a training set $\mathcal{S}_{M}:=\left\{(x_{0},y_{0}),\ldots,(x_{M},y_{M})\right\}\in(X\times Y)^{M}$ consisting of independently and identically distributed samples from an unknown distribution $\mathcal{P}(X\times Y)$ and a hypothesis class $\mathcal{H}$. The objective is to identify a model $h\in\mathcal{H}$, $h:X\rightarrow Y$, that minimizes the empirical risk
$R_{M}(h):=\frac{1}{M}\sum_{i}L(h(x_{i}),y_{i}),$
@@
$=2(n_{\Phi}(x))^{-2}\sum_{\bar{k}\geq k}\varphi_{\bar{k}}(x)\varphi_{k}(x)\Re\braket{\bar{k}}{U^{\dagger}(\theta)\mathcal{M}_{\theta}U(\theta)}{k}.$ (3)
Note that for the $x$-dependence we obtain products $\varphi_{\bar{k}}(x)\varphi_{k}(x)$. To include the original piecewise polynomials from Section 2 within this functional space, we set $\varphi_{k}:=\sqrt{\hat{\varphi}_{k}}$. Due to property (III) this automatically normalizes the basis vector so that $n_{\Phi}(x)=1$. The entire set of functions $\varphi_{\bar{k}}\varphi_{k}$ is plotted in Figure 4.
@@
To illustrate this point, consider Figure 5. These are plots of $h(x,\theta=0)$ for $x\in[-1,1]$ and different observables $\mathcal{M}$, i.e., the data embedding contains no trainable layers and only parameter-independent observables. From (5.2), we see that observables diagonal in the computational $Z$-basis correspond to combinations of piecewise polynomials $\hat{\varphi}_{k}$ only. The qubit number on which we measure reflects the frequency of sign changes. Vice versa, purely off-diagonal observables, like the Pauli-$X$, correspond to non-polynomial elements $\sqrt{\hat{\varphi}_{\bar{k}}\hat{\varphi}_{k}}$, $\bar{k}\neq k$. Similar statements hold for products of these observables.
Figure 5: Plots of only PPTNQFE and different observables, no variational layers or trainable parameters.
In Figure 6, we approximate a discontinuous jump function with added Gaussian noise. Traditional rotation-based embeddings will struggle with such functions, necessitating high frequencies to capture discontinuities. On the contrary, PPTNQFE can achieve a satisfactory approximation with two trainable parameters. For PPTNQFE, we use no variational layers and a parameter-dependent observable $\mathcal{M}_{\theta}=\theta_{0}I+\theta_{1}Z_{0}$. For the rotation-based model, we use exponential feature encoding as in Figure 2, the same observable and 3 variational layers, resulting in a total of 89 trainable parameters.
@@
  • a single variational layer consisting of universal single-qubit gates and CNOTs, combined with 5 trainable observables $\theta_{k}Z_{k}$, and a constant bias term, culminating in a total of 45 trainable parameters;
@@
For PPTNQFE, we use an MPS-structured variational layer, similar to Figure 2, but with three 2-qubit gates instead of a single layer. The rationale for employing an MPS-structured variational layer is based on the idea that the function $\sin$ can be represented as a rank-2 MPS, as detailed in Oseledets (2013). By aligning the variational layer with a rank-2 MPS structure, the aim is to discover a unitary transformation that effectively maps back to the computational basis. This method has shown improved outcomes compared to the use of generic variational layers, although it remains a heuristic approach at this stage. We use observables similar to those in the rotation-based model, $\mathcal{M}_{\theta}=\theta_{0}+\sum_{k}\theta_{k+1}Z_{k}$, resulting in 93 trainable parameters in total. The outcomes are presented in Figure 7.
@@
Both rotation-based models overfit noise, with the simpler one performing slightly better. This is mainly due to the accessible spectrum being too large and no bias towards the correct solution. The number of parameters was deliberately taken to be larger than necessary to demonstrate noise overfitting. Similarly, PPTNQFE with 93 parameters does not perform as well as for the jump function, with square root cusps introduced due to the added variational layers in the model.
@@ 6 Conclusions
We explored two applications of this method in one-dimensional settings. First, PPTNQFE allows efficient point evaluations of discretized solutions to PDEs, enhancing the appeal of quantum algorithms for solving PDEs by addressing data read-out challenges. Second, we demonstrated the utility of PPTNQFE in function regression, particularly highlighting its aptitude for modeling functions with localized features, such as jump discontinuities. An interesting open question is how to design problem-specific trainable circuits suited for approximating more complex dependencies with local effects.

### v5

@@ Abstract
This work introduces a novel method for embedding continuous variables into quantum circuits via piecewise polynomial features, utilizing low-rank tensor networks. Our approach, termed Piecewise Polynomial Tensor Network Quantum Feature Encoding (PPTNQFE), aims to broaden the applicability of quantum algorithms by incorporating spatially localized representations suited for numerical applications like partial differential equations and function regression. We demonstrate the potential of PPTNQFE through efficient point evaluations of solutions of discretized differential equations and in modeling functions with localized features such as jump discontinuities. While promising, challenges such as unexplored noise impact and design of trainable circuits remain.
@@ 1.1 Quantum Advantage in Quantum Machine Learning
The rationale for employing PQCs often hinges on the large size of the quantum state space. Nonetheless, these circuits often produce Fourier series models, which may not fully take advantage of the exponentially large Hilbert space and can be emulated classically Schuld et al. (2021); Kordzanganeh et al. (2023); Shin et al. (2023), or they may be overly expressive, hindering effective training and generalization Kübler* et al. (2021); Huang et al. (2021b); Jerbi et al. (2023); Cerezo et al. (2023). Furthermore, the impact of statistical and hardware-induced errors is frequently overlooked, potentially undermining any claims of enhanced expressiveness or precision.
@@ 1.2 Inductive Bias and Feature Encoding
The efficacy of a model extends beyond mere expressiveness; it depends on the synergy between sufficient expressiveness and the appropriate, problem-specific inductive bias. Recent innovative studies have explored the potential to encode specific biases more efficiently using quantum models Bowles et al. (2023); Gili et al. (2023); Nguyen et al. (2022); Schatzki et al. (2024).
@@ 1.3 This Work
Our approach leverages an intermediate tensor network (TN) representation, which makes it adaptable to any function that can be represented in a low-rank format Oseledets (2013); Ali and Nouy (2023). We denote these embeddings as Piecewise Polynomial Tensor Network Quantum Feature Encodings (PPTNQFE). It is crucial to note that these embeddings are not inherently superior or inferior to others and do not confer any QA. Through toy examples, we elucidate the scenarios where PPTNQFEs are advantageous and when they are not.
The structure of this paper is as follows. Section 2 reviews piecewise polynomials and the properties required for embedding. Section 3 discusses TNs and functional TN representations. Section 4 details the methodology for expressing piecewise polynomial bases as low-rank TNs and their subsequent conversion into quantum circuits. Section 5 presents three important applications: point evaluations of PDE solutions, function learning (regression) and classification with quantum kernels. The paper concludes in Section 6. All computations with quantum algorithms were performed using a state vector simulator.
@@ 2 Piecewise Polynomials
A plot of the hat functions is presented in Figure 1. These satisfy conditions (I), (II) and (III). Moreover, by taking tensor products of the proposed feature embeddings, we can trivially extend them to any spatial dimension. See Figure 2 for a two-dimensional example.
Figure 1: 1D piecewise linear hat basis for a grid of 8 uniformly spaced points, including the boundary $x\in\{\pm 1\}$.
Figure 2: 2D bilinear basis functions formed as tensor products of 4 linear basis functions in each coordinate direction, i.e., 16 functions in total.
@@ 3 Tensor Networks
Tensor networks are an implicit, parsimonious representation of high-order tensors. They are most commonly known for their effectiveness in modeling high-dimensional quantum systems with low entanglement. For details on TN formats and diagrams, see Biamonte and Bergholm (2017). In this work, we will be using the matrix product state (MPS) format Schollwöck (2011), also known under the name tensor train (TT) format Oseledets (2011), see Figure 3. A state |$\ket{\psi}$ over $n$ sites is called an MPS (with open boundary conditions) if it can be expressed as
|$\ket{\psi}=\sum_{s_{0},\ldots,s_{n-1}}C^{0}_{s_{0}}\cdots C^{n-1}_{s_{n-1}}\ket{s_{0}\ldots s_{n-1}}.$
@@ 4 Methodology
The case of even $k$ is similar. Since the maximum rank is $2$, converting to a quantum circuit (see Figure 3) we require only $2$-qubit gates.
@@
Finally, having embedded 1D features, we can easily extend this approach to any number of features by taking tensor products of the MPSs. For instance, given three features $x$, $y$, and $z$ with corresponding binary indices $i_{k}$, $j_{k}$, and $m_{k}$, one possible construction is the canonical ordering:
$\Phi(x,y,z)=$ $A^{0}(x)[\cdot,i_{0},\cdot]\cdots A^{n-1}(x)[\cdot,i_{n-1},\cdot]\cdot$
$B^{0}(y)[\cdot,j_{0},\cdot]\cdots B^{n-1}(y)[\cdot,j_{n-1},\cdot]\cdot$
$C^{0}(z)[\cdot,m_{0},\cdot]\cdots C^{n-1}(z)[\cdot,m_{n-1},\cdot].$
Another commonly used alternative is the transposed ordering
$\Phi(x,y,z)=$ $A^{0}(x)[\cdot,i_{0},\cdot]\cdot B^{0}(y)[\cdot,j_{0},\cdot]\cdot C^{0}(z)[\cdot,m_{0},\cdot]\cdots$
$A^{n-1}(x)[\cdot,i_{n-1},\cdot]\cdot B^{n-1}(y)[\cdot,j_{n-1},\cdot]\cdot C^{n-1}(z)[\cdot,m_{n-1},\cdot].$
For further details on the effects of different orderings, see Kazeev et al. (2017); Markeeva et al. (2021); Ali and Nouy (2023). All orderings yield a feature vector $\Phi(x,y,z)$ with the same values but arranged in different sequences.
Remark 4.2
We emphasize that constructing the MPS for this feature encoding is computationally efficient, with a complexity of $\mathcal{O}(r)$. Specifically, in the example discussed in this paper, where $r=2$, the construction is particularly inexpensive.
@@ 5 Applications
In this section, we dive into three illustrative applications of the PPTNQFE method that we have introduced. The first application focuses on the efficient point evaluation of solutions to differential equations. In this scenario, the discretized and normalized solution is encoded in a quantum state, with the basis coefficients represented within the quantum amplitudes. The second application we explore is regression, specifically the supervised learning of functions depending on continuous variables from given data samples. Here, PPTNQFE allows, e.g., the approximation of jump functions in a particularly simple manner. Finally, the third application is classification with quantum kernels. We demonstrate on synthetic datasets that, in some cases, PPTNQFE achieves similar results to rotation-based encoding. In other cases, PPTNQFE can avoid misclassification due to the periodic nature of rotation-based encodings.
@@ 5.1 Differential Equations
The associated (rescaled) linear system, represented in matrix form,
$\boldsymbol{A}\boldsymbol{u}=\begin{pmatrix}2&-1&0&\dots&\\ -1&2&-1&\dots&\\ &&\ddots&&\\ &&&&-1\\ &&&-1&2\end{pmatrix}\boldsymbol{u}=\boldsymbol{f}.$
@@ 5.2 Regression
In the context of supervised learning, we are presented with a training set $\mathcal{S}_{M}:=\left\{(x_{0},y_{0}),\ldots,(x_{M},y_{M})\right\}\in(X\times Y)^{M}$ consisting of independently and identically distributed samples from an unknown distribution $\mathcal{P}(X\times Y)$ and a hypothesis class $\mathcal{H}$. The objective is to identify a model $h\in\mathcal{H}$, $h:X\times\Theta\rightarrow Y$, that minimizes the empirical risk
$R_{M}(h):=\frac{1}{M}\sum_{i}L(h(x_{i},\theta),y_{i}),$
@@
$=2(n_{\Phi}(x))^{-2}\sum_{\bar{k}\geq k}\varphi_{\bar{k}}(x)\varphi_{k}(x)\Re\braket{\bar{k}}{U^{\dagger}(\theta)\mathcal{M}_{\theta}U(\theta)}{k},$ (3)
for any unitary $U(\theta)$. Note that for the $x$-dependence we obtain products $\varphi_{\bar{k}}(x)\varphi_{k}(x)$. To include the original piecewise polynomials from Section 2 within this functional space, we set $\varphi_{k}:=\sqrt{\hat{\varphi}_{k}}$. Due to property (III) this automatically normalizes the basis vector so that $n_{\Phi}(x)=1$. The entire set of functions $\varphi_{\bar{k}}\varphi_{k}$ is plotted in Figure 5.
@@
To illustrate this point, consider Figure 6. These are plots of $h(x,\theta=0)$ for $x\in[-1,1]$ and different observables $\mathcal{M}$, i.e., the data embedding contains no trainable layers and only parameter-independent observables. From (5.2), we see that expectations values of operators that are diagonal in the computational $Z$-basis correspond to combinations of piecewise polynomials $\hat{\varphi}_{k}$ only. The qubit number on which we measure reflects the frequency of sign changes. Vice versa, purely off-diagonal observables, like the Pauli-$X$, correspond to non-polynomial elements $\sqrt{\hat{\varphi}_{\bar{k}}\hat{\varphi}_{k}}$, $\bar{k}\neq k$. Similar statements hold for products of these observables.
Figure 6: Plots of expectation values of different observables, applied immediately after the PPTNQFE embedding – no variational layers or trainable parameters.
The following training of quantum regression models was performed on quantum state vector simulator with AMSGrad, using the MSE loss on a training set of 50 equally spaced points. More details can be found in the notebooks, see Section 7.
In Figure 7, we approximate a discontinuous jump function with added Gaussian noise. Traditional rotation-based embeddings will struggle with such functions, necessitating high frequencies to capture discontinuities. On the contrary, PPTNQFE can achieve a satisfactory approximation with two trainable parameters. For PPTNQFE, we use no variational layers and a parameter-dependent observable $\mathcal{M}_{\theta}=\theta_{0}I+\theta_{1}Z_{0}$, where $Z_{0}$ is the $Z$-operator applied to the 00’th qubit. For the rotation-based model, we use exponential feature encoding as in Figure 3, the same observable and 3 variational layers, resulting in a total of 89 trainable parameters.
@@
  • a single variational layer consisting of universal single-qubit gates and CNOTs, combined with 5 trainable observables $\theta_{k}Z_{k}$ ($Z_{k}$ is the $Z$-operator applied to qubit $k$), and a constant bias term, culminating in a total of 45 trainable parameters;
@@
For PPTNQFE, we use an MPS-structured variational layer, similar to Figure 3, but with three 2-qubit gates instead of a single layer. The rationale for employing an MPS-structured variational layer is based on the idea that the $\sin$ function can be represented as a rank-2 MPS, as detailed in Oseledets (2013). By using a variational layer with a similar structure to the PPTNQFE embedding layer, the aim is to discover a unitary transformation that effectively maps back to the computational basis. This method has shown improved outcomes compared to the use of generic variational layers, although it remains a heuristic approach at this stage. We use observables similar to those in the rotation-based model, $\mathcal{M}_{\theta}=\theta_{0}+\sum_{k}\theta_{k+1}Z_{k}$, resulting in 93 trainable parameters in total. The outcomes are presented in Figure 8.
@@
Both rotation-based models overfit noise, with the simpler one performing slightly better. This is mainly due to the accessible spectrum being too large and no bias towards the correct solution. The number of parameters was deliberately taken to be larger than necessary to demonstrate noise overfitting. Similarly, PPTNQFE with 93 parameters does not perform as well as for the jump function, with square root cusps introduced due to the added variational layers in the model. The occurrence of non-polynomial terms was explained in Figure 6.
@@
5.3 Classification
For the final example, we consider classification with quantum kernels. As pointed out in Schuld and Petruccione (2021), supervised QML methods without data re-uploading are kernel methods. Thus, we can determine the optimal solution by minimizing the hinge loss over a training set through inverting the kernel matrix. If $x\mapsto\rho(x)$ is the embedding map, the kernel map is given by
$\kappa(x,x^{\prime})=\operatorname{Tr}[\rho(x^{\prime})\rho(x)],$
and the prediction for any new point $x^{*}$ is computed as
$f(x^{*})=\sum_{m=1}^{M}\alpha_{m}\kappa(x^{*},x_{m}),$
where $M$ is the number of training samples and the coefficients $\alpha_{m}$ are determined during training.
We apply this method to both the angle-based embedding and PPTNQFE, and test the quantum kernel classification on two synthetic datasets, as shown in Figures 10 and 11. For a fair comparison, we use 6 qubits for both models. For the angle-based embedding, this corresponds to a Fourier spectrum of 13 different frequencies per feature (169 in total), and for PPTNQFE, this corresponds to 8 hat functions for each feature (64 in total).
Both models perform similarly well on the two half moons dataset. Some points are misclassified due to noise, which places some points of different classes very close to each other.
For the line-like clusters in Figure 11, a simple linear model should already perform well. However, the angle-based embedding misclassifies some points that are far from the clusters due to the periodic nature of the embedding and the presence of high-frequency components in the model. On the other hand, PPTNQFE does not suffer from this issue and only misclassifies one point that is close to the boundary between the two clusters.
Figure 10: Top left: full dataset of two half moons representing two classes with noise. The training set for both quantum kernels was chosen as 80% of the full dataset, randomly selected. Top right: classification results on the full dataset for the angle-based embedding after training. Bottom: classification results on the full dataset for PPTNQFE after training.
Figure 11: Top left: full dataset of three lines representing two classes with noise. The training set for both quantum kernels was chosen as 80% of the full dataset, randomly selected. Top right: classification results on the full dataset for the angle-based embedding after training. Bottom: classification results on the full dataset for PPTNQFE after training.
@@ 6 Conclusions
We explored three applications of this method. First, PPTNQFE allows efficient point evaluations of discretized solutions to PDEs, enhancing the appeal of quantum algorithms for solving PDEs by addressing data read-out challenges. Second, we demonstrated the utility of PPTNQFE in function regression, particularly highlighting its aptitude for modeling functions with localized features, such as jump discontinuities. Third, we showed that PPTNQFE can easily identify non-linear decision boundaries, similar to angle-based embeddings, while avoiding some of the typical issues associated with the periodic Fourier nature of angle-based embeddings. An interesting open question is how to design problem-specific trainable circuits suited for approximating more complex dependencies with local effects.